OpenStudy (anonymous):

Let f be the function with derivative : sin(x^2+1). How many relative extrema does f have on the interval 2

OpenStudy (anonymous):

Set the derivative equal to zero and solve for x: 0 = sin(x² + 1) kπ = x² + 1 kπ - 1 = x² ±√(kπ - 1) = x K is an integer, so plug in values for k and solve for x-values in between 2 and 4. When k = 2, 3, 4, and 5 it works out, thus: x = √(2π - 1), √(3π - 1), √(4π - 1), and √(5π - 1) are critical points

OpenStudy (anonymous):

Btw I googled it.

OpenStudy (anonymous):

I saw this explanation too but I did not understand the k * pi idea

OpenStudy (anonymous):

sin(x² + 1) = 0 x² + 1 = 2πn or x² + 1 = π(2n+1) x = ± √(2πn - 1) or x = ± √(π(2n+1) - 1) n = 1 or 2 on the interval (2,4) Answer: .. 4 relative extrema This one another answer on google..

OpenStudy (anonymous):

I would surely explain you this if I was Maths student ..>.>

OpenStudy (anonymous):

well thanks for the help. I can figure it out